Registration of all UK Domains (Resolved)
  • Priority - Medium
  • Affecting Other - Domain Registration Services
  • At 07:00 (BST) on Wednesday 10 June 2015 we will have a systems outage lasting about two hours when we will be deploying a software upgrade. During this time all registration requests for .uk, .co.uk, .org.uk and all other TLD's ending in .uk will not be processed immediately. As soon as the software upgrade is complete, we will begin to register all .uk domains requested.We advise not ordering a .uk domain during this time as we will be unable to guarantee registration. After the outage, all registrations made during this period will be processed in order of reciept.We apologise for any inconvenience this may cause you.Technical staff member watching this issue: Peter Bishop. (cPanel Certified Level 1 Technician)

  • Date - 10/06/2015 07:00 - 10/06/2015 15:31
  • Last Updated - 10/06/2015 15:31
Website Upgrade (Resolved)
  • Priority - Critical
  • Affecting Other - Website
  • Please be advised our website will be unavailable on Tuesday 4th November 2014 from approximately 9am while we update & refresh our website, this proces may take up to 6pm. During this time customers will be unable to access all parts of our website including the customer portal, ability to place orders or updates & downgrades.

    We apologise in advance for any inconveniance this may cause.

  • Date - 04/11/2014 09:00 - 09/11/2014 20:56
  • Last Updated - 27/10/2014 22:36
Server Outage (Resolved)
  • Priority - Critical
  • At current, all VPS services on the Skyla node are currently experiencing downtime due to a server outage. We are working hard to resolve this issue and will update you as more information becomes available.

    UPDATE 03/05/2013 21:54: The issue has now been found and is being worked on by our Senior Administrator. Update to follow. (Update by: Peter Bishop)

    UPDATE 03/05/2013 22:52: The issue has now been resolved.

  • Date - 03/05/2013 20:41
  • Last Updated - 03/05/2013 21:52
DNS Issue (Resolved)
  • Priority - Critical
  • Customers on our Kanga server are experiencing an outage due to DNS routing issues with primary and slave nameservers, the latest update to dns records appears to be faulty and has stop the service working correctly.

    Engineers are working on the problem, we aim to have the problem solved within 3 hours.

     

    Update: Fault cleared, a customer had tried setting custom DNS records which were not correct and caused the the service to shutdown.

  • Date - 03/07/2010 11:46 - 03/07/2010 12:04
  • Last Updated - 03/07/2010 12:04
National Datastream Outage (Resolved)
  • Priority - Critical
  • Affecting System - Connectivity & Telephone Systems
  • We have received the following communication from BT Openreach regarding a National Datastream service outage affecting upto 37,500 circuits.

     

    Description of Outage:

    Several BBLCs terminating to Reading and Harbour Exchange are down.

    All datastream customers terminating on rdg-0-dsl, he-1-dsl, and he-2-dsl, are affected by this outage.

     

     

    Geographic location of affected services: (where possible to define).

    Due to the nature of our network affected exchanges could be anywhere in the UK

     

    Customers may experience issues with their ADSL connection, the router will be able to sync and obtain a connection , yet traffic will not pass (Web/Email/VoIP services will be unavailable).

    Updates when received from BT Openreach will be posted here.

    ** 12:17 : 437 exchanges have been confirmed as affected by this outage, Customer Services have received a document detailing affected exchanges.

    ** 12:30 : BT have taken the decision to close North Paddington Exchange due to flooding , LLU services provided by Tiscali at this exchange only are experiencing the same issues as above.

    ***Update 13:45***

    A full list of the affected exchanges can be found here - http://bit.ly/bedra6

    ** 17:30 : The first of 2 generators that BT have ordered for North Paddington Exchange has arrived onsite, once flood water has been cleared engineers will be tasked to install it.

    ***Update 20:25***

    We have had an update from BT. They are saying that it may take up to 4 days to resolve the fault. They will be able to give a more definite time frame to fix once they are allowed inside the site by the fire department.

  • Date - 31/03/2010 06:02 - 08/04/2010 16:53
  • Last Updated - 31/03/2010 23:04
DDOS Attack (Resolved)
  • Priority - Critical
  • SERVER OUTAGE - OWL

    Customers hosted on Owl Server are currently experiencing downtime to due a DDOS Attack.

    Engineers are currently looking into the problem and will get services back online as quickly as possible.

     

    Update at 1926

    Services all running normally and will be monitored for the next 48 hours. Server Downtime of 24 minutes.

  • Date - 01/02/2010 19:10 - 03/02/2010 00:00
  • Last Updated - 03/02/2010 15:14
Email - Unable to Send (Resolved)
  • Priority - Medium
  • Due to a current fault with email service for customers located on server piglet, customers are currently unable to send any mail via our servers, incoming mail appears to be unaffected at this current time.

    Techs are currently looking into the fault. If you need to send email please use your ISPs SMTP server.

     

    Update at 02:27 on Sunday 24th Jan 10:

    Appears a local security update applied to our servers has amended the way customers authenticate before sending email via our servers. Customers need to alter the outgoing mail setting for each email account for which email is sent via our servers, please select log on using same settings as my incoming mail or alternatively enter the same credentials used to retrieve your email.

     

    Customers who have tried sending email in the past hour may not be able to send email for up to 4 hours due to auto spam protection in place which will block sending email if you have failed sending email more than 5 times. If you still cannot send email after 6 hours please contact us.

     

     

  • Date - 24/01/2010 02:15 - 24/01/2010 02:31
  • Last Updated - 24/01/2010 02:31
Customer Portal Upgrade (Resolved)
  • Priority - Critical
  • Affecting System - Customer Portal
  • We have been implementing new customer billing and support systems, this had led to downtime which is now resolved.

    We have reset all passwords on all accounts, please check your email address on file for details, if you have not received this please contact support.

    Thanks

  • Date - 14/01/2010 00:00 - 23/01/2010 18:17
  • Last Updated - 23/01/2010 18:16
My Tudor Net (Resolved)
  • Priority - Critical
  • Affecting System - Customer Billing & Support Platform
  • To enable us to provide a better service we will be upgrading our billing and customer support platform. We aim to have full service restored within 48 hours.

     

    Thanks

    Matthew Taylor

    Network Administrator

  • Date - 12/01/2010 00:00 - 23/01/2010 18:17
  • Last Updated - 11/01/2010 19:07

Server Status

Below is a real-time overview of our servers where you can check if there's any known issues.

Server Name Online Server Load Uptime
Drayden
Hypno
PBX
Rudy
Snover
Thunder
Voltage

Powered by WHMCompleteSolution